OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, find the limit. (If an answer does not exist, enter DNE.) an = ln(n + 4) -ln n

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

write it as \[\ln(\frac{n+4}{n})\] then you see it converges to \(\ln(1)=0\)

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

For a more rigorous approach, you might want to establish that the sequence is bounded and monotonic. Any bounded/monotonic sequence must converge (see here: ` http://en.wikipedia.org/wiki/Monotone_convergence_theorem `).